# Defensive Prompt Patch: A Robust and Interpretable Defense of LLMs against Jailbreak Attacks

### v1

@@ Abstract
Safety, security, and compliance are essential requirements when aligning large language models (LLMs). However, many seemingly aligned LLMs are soon shown to be susceptible to jailbreak attacks. These attacks aim to circumvent the models’ safety guardrails and security mechanisms by introducing jailbreak prompts into malicious queries. In response to these challenges, this paper introduces Defensive Prompt Patch (DPP), a novel prompt-based defense mechanism specifically designed to protect LLMs against such sophisticated jailbreak strategies. Unlike previous approaches, which have often compromised the utility of the model for the sake of safety, DPP is designed to achieve a minimal Attack Success Rate (ASR) while preserving the high utility of LLMs. Our method uses strategically designed interpretable suffix prompts that effectively thwart a wide range of standard and adaptive jailbreak techniques. Empirical results conducted on LLAMA-2-7B-Chat and Mistral-7B-Instruct-v0.2 models demonstrate the robustness and adaptability of DPP, showing significant reductions in ASR with negligible impact on utility. Our approach not only outperforms existing defense strategies in balancing safety and functionality, but also provides a scalable and interpretable solution applicable to various LLM platforms.
@@ 1 Introduction
Recent advances in large language models (LLMs) [25, 31] such as GPT-4 [18], LLAMA-2 [2], and Mistral [7] have showcased their ability to understand and generate text akin to human interaction [26, 27, 32]. These models, powered by the Transformer architecture, excel in processing sequential data and understanding complex language patterns, hence enhancing tasks like text summarization, creative writing, and coding. To maintain model integrity and mitigate undesired outputs, developers implement alignment constraints using techniques like Reinforcement Learning with Human Feedback (RLHF) [22, 23, 24] and Supervised Fine-Tuning (SFT).
Despite these alignment efforts, current LLMs can be tricked to generate undesirable output, as demonstrated by various jailbreak attacks [1, 3, 5, 4]. Initial strategies like the GCG attack involve crafting adversarial suffixes combined with user queries to manipulate model outputs [1]. More sophisticated techniques such as the AutoDAN [3], PAIR [5], and TAP [4] attacks generate interpretable jailbreak templates that enhance the efficacy and readability of the attacks.
In response to these vulnerabilities, the development of defensive strategies [19, 20, 28] has become increasingly vital. Prompt-based defenses, such as Self-Reminder [10], Goal Prioritization [11], and RPO [12], involve improving system prompts to enhance LLM alignment. These methods are simple yet effective, requiring minimal in-depth model knowledge and sparing model re-training as they operate at the text input level.
Nevertheless, these prompt-based defense mechanisms frequently grapple with the trade-off between preserving utility and effectively mitigating jailbreaks. Although Goal Prioritization excels in defense, it substantially compromises model utility. On the other hand, RPO retains utility but provides limited defense coverage. While Self-Reminder achieves a better balance, it fails to deliver satisfactory performance on more aligned models such as LLAMA-2-7B-Chat, owing to deficiencies in its search algorithm for the optimal prompt. We provide a comparative analysis of different prompt-based defenses in Table 1.
Optimizable Prompt Gradient-Based Search Interpretable Attack Success Rate Utility Degradation
@@
Figure 1: Overview of Defensive Prompt Patch. (a) showcases an example of jailbreak attacks. (b) is the DPP training phase in which the algorithm takes in the refusal and helpful datasets and a prototype of the defense prompt. Then, the algorithm forms the defense prompt population by revising the prototype using LLM. For each of the defense prompts in the population, the algorithm will evaluate the defense and utility scores as detailed in Sec. 3. The algorithm keeps editing the defense prompts with low scores using the Hierarchical Genetic Search algorithm. (c) shows the deployment of DPP in the LLM inference phase, by adding the best DPP in (b) (indicated in green patch) to every input query. (d) shows the trade-off graphs between the win-rate (utility) [14] and attack success rate (ASR) in both LLAMA-2-7B-Chat and Mistral-7B-Instruct-v0.2 models for different defenses.
To address these deficiencies, we introduce Defensive Prompt Patch (DPP), a novel, prompt-based defense mechanism. As illustrated in Figure 1, DPP uses adversarial and utility datasets to iteratively optimize and refine a suffix prompt to be appended to every input query for balancing alignment and utility. Figure 1(d) demonstrates that DPP notably reduces the Attack Success Rate (ASR) to 3.8% on the LLAMA-2-7B-Chat model without compromising utility. Furthermore, it extends robust defense capabilities to less-aligned models, such as the Mistral-7B-Instruct-v0.2, where it achieves a significant reduction in ASR to 2.0% while maintaining minimal utility loss.
@@
  • Improved Defense with Minimal Utility Trade-off: DPP is designed to minimize jailbreak risks while maintaining high utility, addressing the common pitfalls in current prompt-based defenses. Figure 1(d) summarizes its superior performance in balancing jailbreak risk and utility (Win-Rate).
@@
  • Interpretability and Stability of Prompt-based Defenses: We examined the best DPP found by our algorithm and demonstrated its enhanced interpretability over existing prompt-based defenses. We also conducted an ablation study on the LLAMA-2-7B-Chat model to validate that using DPP as a suffix to every input query attains better defense and utility compared with using it as a prefix.
@@ 2 Related Work
Jailbreak attacks have evolved through various innovative mechanisms. For instance, techniques like the PAIR and TAP Attacks [5, 4] automate the creation of jailbreak prompts using a secondary “attacker” LLM, which poses serious threats through black-box access to the target LLM. Similarly, the ICA Attack [8] leverages in-context learning to misaligned responses, and the Catastrophic Attack [9] manipulates generation configurations to trigger misaligned outputs. GCG Attack [1] optimize adversarial inputs using gradient-based approaches, and the AutoDAN Attack [3] employs genetic algorithms to refine prompts based on specific templates. Another notable method, the Base64 Attack [6], encodes malicious queries in Base64 to bypass content filters subtly.
Defensive strategies have been developed in response to these sophisticated attacks to reinforce the security of LLMs. Techniques such as the Self-Reminder [10] defense modify the system prompt of LLMs to induce more self-aware and aligned processing. The RPO (Robust Prompt Optimization) [12] modifies objectives to minimize the perceptual distance between harmful queries and safe responses. Furthermore, Goal Prioritization and Default System Prompts [11, 15] are designed to direct LLMs to prioritize safety and prevent the generation of harmful outputs.
These attacks and defenses represent a dynamic interplay between the capabilities of LLMs and the measures required to secure them. Detailed descriptions and evaluations of these defense methods will be further discussed in the Sec. 4 section, where their effectiveness against various adversarial strategies is systematically analyzed.
@@ 3.1 Preliminaries
Jailbreak Attack: A jailbreak attack on an LLM aims to circumvent model alignment by using meticulously crafted prompts [29, 30]. We denote a malicious query as $\mathbf{u}_{1:n}=\langle u_{1},u_{2},\dots,u_{n}\rangle$, with each $u_{i}$ being an input token. Ordinarily, the LLM would reject such queries based on its alignment policies. However, refined jailbreak queries, $\tilde{\mathbf{u}}_{1:m}=\langle\tilde{u}_{1},\tilde{u}_{2},\dots,\tilde{u}_{m}\rangle$, manipulate these policies to elicit a compliant response $\mathbf{r}_{1:k}=\langle r_{1},r_{2},\dots,r_{k}\rangle$, reflecting the original malicious intent.
@@
Utility Degradation: We measure utility degradation by the deviation in LLM responses to benign queries appended with $\mathbf{d}_{1:l}$. Ideally, the response to a benign query $\mathbf{b}_{1:p}=\langle b_{1},b_{2},\dots,b_{p}\rangle$ patched by $\mathbf{d}_{1:l}$ should closely match the response to $\mathbf{b}_{1:p}$ alone.
Mathematical Formulation: We define the $\oplus$ operation as the concatenation of two sequences. For a given sequence $\mathbf{a}_{1:n}=\langle a_{1},\dots,a_{n}\rangle$ and $\mathbf{z}_{1:m}=\langle z_{1},\dots,z_{m}\rangle$, $\mathbf{a}_{1:n}\oplus\mathbf{z}_{1:m}$ is defined as: $\mathbf{a}_{1:n}\oplus\mathbf{z}_{1:m}=\langle a_{1},\dots a_{n},z_{1},\dots z_{m}\rangle$. We denote sequences of harmful responses and jailbreak inputs by $\mathbf{r}_{1:k}$ and $\tilde{\mathbf{u}}_{1:m}$, respectively. Since LLMs are specifically trained to predict the probability of the next word, we define the goal (i.e., the objective function to be maximized) of a jailbreak attack as:
@@
where $\mathbf{h}_{1:q}$ is the normal response for each benign queries $\mathbf{b}_{1:p}$. The DPP algorithm’s efficacy is evaluated by its performance in both defense against malicious queries and impact on utility on benign queries.
@@ 3.2 Score Evaluation
where ${S}_{D_{i}}$ denotes the refusal score attributed to the $i$-th DPP within the population of DPPs. The vector $\mathbf{s}_{1:n}$ encapsulates the refusal response, $\tilde{\mathbf{u}}_{1:m}$ represents the jailbreak query, and $\mathbf{d}_{1:l}$ is the our defensive mechanism.
@@
where ${S}_{H_{i}}$ represents the helpfulness score assigned to the $i$-th DPP within the population of DPPs. The vector $\mathbf{h}_{1:q}$ denotes the standard response, whereas $\mathbf{b}_{1:p}$ refers to the benign query. The overall score function for training DPP combines the refusal and helpful scores, weighted by coefficients $\alpha$ and $\beta$, respectively:
@@ 3.3 DPP Training Algorithm
Using the total score defined in Sec. 3.2, we use a Hierarchical Genetic Algorithm (HGA) to optimize DPP, drawing inspiration from the AutoDAN jailbreak attack in [3]. We adapt and extend HGA to iteratively refine DPP based on our defined scores, as depicted in Figure. 1 (b) and (c). to develop our methodology, which we term the Defensive Prompt Patch Algorithm (DPP Algorithm).
Initially, we establish a baseline DPP, designated as the prototype. Without loss of generality, this prototype may take the form of either a Prefix DPP or a Suffix DPP. The relative effectiveness of each configuration is assessed in Appendix. D. Following this, the prototype is subjected to $K$ iterations of rewriting via an LLM to potentially refine the DPP, creating a population of DPP candidates. Each candidate within the population is evaluated by sampling refusal data pairs and helpful data pairs from adversarial/utility datasets to compute the total score, as formulated in Eq. 6. Details on adversarial/utility datasets in our implementation can be found in Sec. 4.1.
@@
These processes—Sentence-Level Word Substitution and Paragraph-Level Sentence Swap and Mutations—aim to increase the diversity within the defensive prompt patch population and enhance the likelihood of identifying the optimal patch.
The full algorithm is delineated in Algorithm 1. Ultimately, the algorithm produces an updated set of optimized DPPs, comprising $K$ enhanced patches, and identifies the Best Defensive Prompt Patch based on the highest total score.
@@
1:Arguments: Defensive Prompt Patch Prototype $O$ , refusal pair $(x^{r},y^{r})$, helpful pair $(x^{h},y^{h})$, $\alpha$ and $\beta$, target LLM
2:Initialization: Number of optimization iteration $I$, batch size, $p_{crossover}$, $p_{mutate}$, Sentence-level iterations, Paragraph-level iterations, number of steps, number of parent set size
@@
4:while $I$ is not reached do
5:     for $iteration$ in sentence-level iterations do
6:         Evaluate refusal/helpful score of each DPP with $(x^{r},y^{r})$/$(x^{h},y^{h})$ and target LLM
7:         Final Score $\leftarrow$ calculate the score using Eq. (6)
8:         Select elite and parent prompts from DPP_Set according to Final Score

### v2

# Defensive Prompt Patch: A Robust and Generalizable Defense of Large Language Models against Jailbreak Attacks

@@ Abstract
Safety, security, and compliance are essential requirements when aligning large language models (LLMs). However, many seemingly aligned LLMs are soon shown to be susceptible to jailbreak attacks. These attacks aim to circumvent the models’ safety guardrails and security mechanisms by introducing jailbreak prompts into malicious queries. In response to these challenges, this paper introduces Defensive Prompt Patch (DPP) [^1] Link to DPP Project Page and Code: https://huggingface.co/spaces/TrustSafeAI/Defensive-Prompt-Patch-Jailbreak-Defense, a novel prompt-based defense mechanism specifically designed to protect LLMs against such sophisticated jailbreak strategies. Unlike previous approaches, which have often compromised the utility of the model for the sake of safety, DPP is designed to achieve a minimal Attack Success Rate (ASR) while preserving the high utility of LLMs. Our method uses strategically designed suffix prompts that effectively thwart a wide range of standard and adaptive jailbreak techniques. Empirical results conducted on Llama-2-7B-Chat and Mistral-7B-Instruct-v0.2 demonstrate the robustness and adaptability of DPP, showing significant reductions in ASR with negligible impact on utility. Our approach not only outperforms existing defense strategies in balancing safety and functionality, but also provides a scalable and robust solution to various LLM platforms.
Defensive Prompt Patch: A Robust and Generalizable Defense of
Large Language Models against Jailbreak Attacks
Chen Xiong1, Xiangyu Qi2, Pin-Yu Chen3, Tsung-Yi Ho1 1Department of Computer Science and Engineering, The Chinese University of Hong Kong 2Princeton University  3IBM Research Correspondence: {cxiong23, tyho}@cse.cuhk.edu.hk, xiangyuqi@princeton.edu, pin-yu.chen@ibm.com
@@ 1 Introduction
Recent advances in large language models (LLMs) (Vaswani et al., 2023; Devlin et al., 2019) such as GPT-4 (OpenAI, 2023), Llama-2 (Touvron et al., 2023), and Mistral (Jiang et al., 2023) have showcased their ability to understand and generate text akin to human interaction (Zhong et al., 2023; Pu et al., 2023; Dasgupta et al., 2023). These models, powered by the Transformer architecture, excel in processing sequential data and understanding complex language patterns, hence enhancing tasks like text summarization, creative writing, and coding. To maintain model integrity and mitigate undesired outputs, developers implement alignment constraints using techniques like Reinforcement Learning with Human Feedback (RLHF) (Askell et al., 2021; Bai et al., 2022; Ouyang et al., 2022) and Supervised Fine-Tuning (SFT) (Zhang et al., 2024a; Tajwar et al., 2024).
Despite these alignment efforts, current LLMs can be tricked to generate undesirable output, as demonstrated by various jailbreak attacks (Zou et al., 2023; Liu et al., 2023; Chao et al., 2023; Mehrotra et al., 2023). Initial strategies like the GCG attack (Zou et al., 2023) involve crafting adversarial suffixes combined with user queries to manipulate model outputs. More sophisticated techniques such as the AutoDAN (Liu et al., 2023), PAIR (Chao et al., 2023), and TAP (Mehrotra et al., 2023) attacks generate interpretable jailbreak templates, improving attack efficacy and readability.
In response to these vulnerabilities, the development of defensive strategies (Jain et al., 2023; Robey et al., 2023; Zhang et al., 2024b) has become increasingly vital. Prompt-based defenses, such as Self-Reminder (Xie et al., 2023), Goal Prioritization (Zhang et al., 2023b), and RPO (Zhou et al., 2024), involve improving system prompts to enhance LLM alignment. These methods demonstrate a balance of simplicity and effectiveness, requiring minimal detailed knowledge of the model architecture. They operate at the text input level, thereby eliminating the need for any additional model re-training.
Nevertheless, these prompt-based defense mechanisms frequently grapple with the trade-off between preserving utility and effectively mitigating jailbreaks. Although Goal Prioritization excels in defense, it substantially compromises model utility. On the other hand, RPO retains utility but provides limited defense coverage. While Self-Reminder achieves a better balance, it fails to deliver satisfactory performance on more aligned models such as Llama-2-7B-Chat, owing to deficiencies in its search algorithm for the optimal prompt. To elucidate these findings, we present a comparative analysis of various prompt-based defense strategies in Table 1.
Optimizable Prompt Logits-Based Search Human Understandability Attack Success Rate Utility Degradation
@@
Figure 1: Overview of Defensive Prompt Patch. (a) showcases an example of jailbreak attacks. (b) is the DPP training phase in which the algorithm takes in the refusal and helpful datasets and a prototype of the defense prompt. Then, the algorithm forms the defense prompt population by revising the prototype using LLM. For each of the defense prompts in the population, the algorithm will evaluate the defense and utility scores as detailed in Sec. 3. The algorithm keeps editing the defense prompts with low scores using the Hierarchical Genetic Search algorithm. (c) shows the deployment of DPP in the LLM inference phase, by adding the best DPP in (b) (indicated in green patch) to every input query. (d) shows the trade-off graphs between the win-rate (utility) (Li et al., 2023) and attack success rate (ASR) in both Llama-2-7B-Chat and Mistral-7B-Instruct-v0.2 for different defenses.
To address these deficiencies, we introduce Defensive Prompt Patch (DPP), a novel, prompt-based and human-readable defense mechanism (see Appendix A for motivation). At its core, DPP is learned with a Hierarchical Genetic Algorithm (HGA). The algorithm alternates between (i) sentence-level word substitutions and (ii) paragraph-level cross-overs, retaining only the prompt patches that raise the model’s refusal likelihood on harmful inputs while preserving helpfulness on benign ones. After a few iterations, the search produces a compact suffix prompt that can be attached to any user query—no model retraining required. As illustrated in Figure 1, DPP uses adversarial and utility datasets to iteratively optimize and refine a suffix prompt to be appended to every input query for balancing alignment and utility. Figure 1(d) demonstrates that DPP notably reduces the Attack Success Rate (ASR) to 3.8% on the Llama-2-7B-Chat model without compromising utility. Furthermore, it extends robust defense capabilities to less-aligned models, such as the Mistral-7B-Instruct-v0.2, where it achieves a significant reduction in ASR to 2.0% while maintaining minimal utility loss.
@@
  • Improved Defense with Minimal Utility Trade-off: DPP is designed to minimize jailbreak risks while maintaining high utility, addressing the common pitfalls in current prompt-based defenses. Figure 1(d) summarizes its superior performance in balancing jailbreak risk and utility.
@@
  • Clarity and Stability of Prompt-based Defenses: We examined the best DPP found by our algorithm and demonstrated its enhanced clarity over existing prompt-based defenses. In addition, we conducted an ablation study on the Llama-2-7B-Chat model to validate that using DPP as a suffix to every input query attains better defense and utility compared with using it as a prefix. Furthermore, we explored the pivotal roles of both utility and defense scores in optimizing the model’s resilience to attacks, while minimizing any potential degradation in performance.
@@ 2 Related Work
Jailbreak attacks have evolved through various innovative mechanisms. For instance, techniques like the PAIR and TAP Attacks (Chao et al., 2023; Mehrotra et al., 2023) automate the creation of jailbreak prompts using a secondary "attacker" LLM, which poses serious threats through black-box access to the target LLM. Similarly, the ICA Attack (Wei et al., 2023b) leverages in-context learning to misaligned responses, and the Catastrophic Attack (Huang et al., 2023) manipulates generation configurations to trigger misaligned outputs. GCG Attack (Zou et al., 2023) optimize adversarial inputs using gradient-based approaches, and the AutoDAN Attack (Liu et al., 2023) employs genetic algorithms to refine prompts based on specific templates. Another notable method, the Base64 Attack (Wei et al., 2023a), encodes malicious queries in Base64 to bypass content filters subtly.
Defensive strategies have been developed in response to these sophisticated attacks to reinforce the security of LLMs. Techniques such as the Self-Reminder (Xie et al., 2023) defense modify the system prompt of LLMs to induce more self-aware and aligned processing. The RPO (Robust Prompt Optimization) (Zhou et al., 2024) modifies objectives to minimize the perceptual distance between harmful queries and safe responses. Furthermore, Goal Prioritization and Default System Prompts (Zhang et al., 2023b; Zheng et al., 2024a, 2023) are designed to direct LLMs to prioritize safety and prevent the generation of harmful outputs.
These attacks and defenses represent a dynamic interplay between the capabilities of large language models (LLMs) and the measures required to secure them. In Section 4, we will provide comprehensive descriptions and evaluations of these defense mechanisms. This section will systematically analyze their effectiveness against a range of adversarial strategies.
@@ 3.1 Preliminaries
Jailbreak Attack: A jailbreak attack on an LLM aims to circumvent model alignment by using meticulously crafted prompts (Yong et al., 2024; Zhang et al., 2023a). We denote a malicious query as $\mathbf{u}_{1:n}=\langle u_{1},u_{2},\dots,u_{n}\rangle$, with each $u_{i}$ being an input token. Ordinarily, the LLM would reject such queries based on its alignment policies. However, refined jailbreak queries, $\tilde{\mathbf{u}}_{1:m}=\langle\tilde{u}_{1},\tilde{u}_{2},\dots,\tilde{u}_{m}\rangle$, manipulate these policies to elicit a compliant response $\mathbf{r}_{1:k}=\langle r_{1},r_{2},\dots,r_{k}\rangle$ that align with the malicious intent, thereby achieving the objectives.
@@
Utility Degradation: We measure utility degradation by the deviation in LLM responses to benign queries appended with $\mathbf{d}_{1:l}$. Ideally, the response to a benign query $\mathbf{b}_{1:p}=\langle b_{1},b_{2},\dots,b_{p}\rangle$ patched by $\mathbf{d}_{1:l}$ should closely match the response to $\mathbf{b}_{1:p}$.
Mathematical Formulation: We define the $\oplus$ operation as the concatenation of two sequences. For two given dummy sequences $\mathbf{a}_{1:n}=\langle a_{1},\dots,a_{n}\rangle$ and $\mathbf{z}_{1:m}=\langle z_{1},\dots,z_{m}\rangle$, $\mathbf{a}_{1:n}\oplus\mathbf{z}_{1:m}$ is defined as: $\mathbf{a}_{1:n}\oplus\mathbf{z}_{1:m}=\langle a_{1},\dots a_{n},z_{1},\dots z_{m}\rangle$. We denote sequences of harmful responses and jailbreak inputs by $\mathbf{r}_{1:k}$ and $\tilde{\mathbf{u}}_{1:m}$, respectively. Since LLMs are specifically trained to predict the probability of the next word, we define the jailbreak goal (i.e., the objective function to be maximized) as:
@@
where $\mathbf{h}_{1:q}$ is the normal response for each benign queries $\mathbf{b}_{1:p}$.
The overall DPP algorithm’s efficacy is evaluated by its performance in both defense against malicious queries and impact on the benign queries.
@@ 3.2 Score Evaluation
where ${S}_{D_{i}}$ denotes the refusal score attributed to the $i$-th DPP within the population of DPPs. The vector $\mathbf{s}_{1:n}$ represents the refusal response, $\tilde{\mathbf{u}}_{1:m}$ represents the jailbreak query, and $\mathbf{d}_{1:l}$ is our DPP.
@@
where ${S}_{H_{i}}$ represents the helpfulness score assigned to the $i$-th DPP within the population of DPPs. The vector $\mathbf{h}_{1:q}$ denotes the standard response, whereas $\mathbf{b}_{1:p}$ refers to the benign query. The overall score function for training DPP combines the refusal and helpful scores. These scores are weighted by the coefficients $\alpha$ and $\beta$, respectively, to balance their contributions within the training process:
@@ 3.3 DPP Training Algorithm
Using the total score defined in Sec. 3.2, we use a Hierarchical Genetic Algorithm (HGA) to optimize DPP, drawing inspiration from the AutoDAN jailbreak attack in (Liu et al., 2023). We adapt and extend HGA to iteratively refine DPP based on our defined scores, as shown in Figure 1 (b) and (c) to develop our methodology, which we call the Defensive Prompt Patch Algorithm (DPP Algorithm). We discuss the key difference between our DPP and AutoDAN in Appendix B.4.
Initially, we establish a baseline DPP, designated as the prototype. Without loss of generality, this prototype may take the form of either a Prefix DPP or a Suffix DPP. The relative effectiveness of each configuration is assessed in Appendix C. Following this, the prototype is subjected to $K$ iterations of rewriting via an LLM to potentially refine the DPP, creating a population of DPP candidates. Each candidate within the population is evaluated by sampling refusal data pairs and helpful data pairs from adversarial/utility datasets to compute the total score, as formulated in Eq. 6. Details on adversarial/utility datasets in our implementation can be found in Sec. 4.1.
@@
These processes—Sentence-Level Word Substitution and Paragraph-Level Sentence Swap and Mutations—aim to increase the diversity within the prompt patch population and enhance the likelihood of identifying the optimal patch.
The full algorithm is delineated in Algorithm 1. Ultimately, the algorithm produces an updated set of optimized DPPs, comprising $K$ enhanced patches, and identifies the Best Defensive Prompt Patch based on the highest total score. A detailed explanation of Algorithm 1 is in Appendix D.
@@
1Arguments: Defensive Prompt Patch Prototype $O$ , refusal pair $(x^{r},y^{r})$, helpful pair $(x^{h},y^{h})$, $\alpha$ and $\beta$, target LLM
2Initialization: Number of optimization iteration $I$, batch size, $p_{crossover}$, $p_{mutate}$, Sentence-level iterations, Paragraph-level iterations, n umber of steps, number of parent set size
@@
4while $I$ is not reached do
5     for $iteration$ in sentence-level iterations do
6         Evaluate refusal/helpful score of each DPP with $(x^{r},y^{r})$/$(x^{h},y^{h})$ and target LLM
7         Final Score $\leftarrow$ calculate the score using Eq. (6)
8         Select elite and parent prompts from DPP_Set according to Final Score